OpenStudy (anonymous):

todd's local bank offers 4% interest compounded monthly. the intermet bank offers 4% compounded daily if todd invests $400 for 10 years what is the difference in the interest earned from the bank? $0.39 $1.49 $3.90 $14.90

OpenStudy (misty1212):

\[400\times (1+\frac{.04}{365})^{3650}-400\times (1+\frac{.04}{12})^{120}\]should do it